Mr. Chris's 2018 income and deductions are listed below. Calculated for the following two cases, Mr. Chris's Net Income for tax purposes and indicate any possible loss carryovers from the current year.

Case A

Employment Income.............................28,000

Income(loss) from Business..................(36,000)

Income from property .............................15,000

Taxable Capital Gains..............................21,000

Allowable Capital Losses ........................(27,000)

Subdivision E Deductions (spousal support)....(11,000)

Case B

Employment Income.......................16,000

Income (loss) from Business ..........(21,000)

Income from property........................22,000

Capital Gains ...................................64,000

Capital Losses ..................................(138,000)

Subdivision E Deductions (spousal support) ......(5,000)
Answer #1

Allowable Business Investment Loss(ABIL) can be claimed against any type of income upto 50% of Business Investment Loss.

ABIL for Case A= 50% of 36000= 18000

ABIL for Case B= 50% of 21000= 10500

Remaining 50% can be carried forward for the future years or can be adjusted against any previous year

Carry forward amount for Case A= 18000

Carry forward amount for Case B= 10500

Allowable Capital Losses can be set off only against taxable capital gains. If losses are greater than gains, then remaining amount of losses can be carried forward to future years.

Case A

Allowable Capital Losses= 21000,

Income from Capital Gains = 21000- 21000= Nil

Carry forward= 27000-21000= 6000

Case B

Allowable Capital Losses= 64000,

Income from Capital Gains = 64000-64000= Nil

Carry forward= 138000-64000= 74000

Net Income:

Case A

Employment income 28000

Allowable Business Loss (18000)

Income from property 15000

Income from capital gains Nil

Gross income 25000

Less: Subdivision E (11000)

Deduction

Net taxable income 14000

Case B

Employment income 16000

Allowable Business Loss (10500)

Income from property 22000

Income from capital gains Nil

Gross Income 27500

Less: Subdivision E (5000)

Deduction

Net taxable income 22500